OpenStudy (ledah):

Simplify the following: (x + 6)(x - 6)

OpenStudy (ledah):

I know this is an easy one. but I'm always making this simplest of things difficult

mathslover (mathslover):

\[(x+6)(x-6)=x^2-6^2 = x^2-36\]

mathslover (mathslover):

If you are doing algebra than i will recommend you to learn all the algebraic identities : here (a+b)(a-b) = a^2-b^2 is applied Best of luck

OpenStudy (aravindg):

yep x^2-36

OpenStudy (precal):